OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). I need help

OpenStudy (kropot72):

The first step is to arrange the equation in the form y = mx + b This results in the following y = -3x + 7 The product of the slope of the rearranged given equation, where the slope m = -3, and the slope of the perpendicular line must equal -1. So if we let the slope of the perpendicular line be m we have m * -3 = -1 Giving m = 1/3 So the equation of the perpendicular line at this stage becomes y = (1/3)x + b where the value of the constant b can be found by substituting the given values of x and y for the point. Can you do that?